ultimate value of 'log[(1+d)(1+2d)(1+3d)...(1+nd)]'

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

A = log[(1+d)(1+2d)(1+3d)...............(1+nd)]

log(ab) = log(a) + log(b)A = log(1+d) + log(1+2d) + log(1+3d) +............... + log(1+nd)A = \sum_{r = 1}^{n}log(1+rd)
A = log(d^{n}(1+\frac{1}{d})_{n})